OpenStudy (briensmarandache):

sin(2x+x)=.5(sqrt(3)) solve

OpenStudy (solomonzelman):

\(\large\color{black}{ \sin(2x+x)=\sin(2x)\cos(x)+\sin(x)\cos(2x)}\) \(\large\color{black}{ \sin(2x+x)=2\sin(x)\cos(x)\cos(x)+\sin(x)(\cos^2x-\sin^2x)}\) \(\large\color{black}{ \sin(2x+x)=2\sin(x)\cos^2(x)+\sin(x)(\cos^2x-\sin^2x)}\) \(\large\color{black}{ \sin(2x+x)=2\sin(x)\cos^2(x)+\sin(x)\cos^2x-\sin^3x}\) \(\large\color{black}{ \sin(2x+x)=3\sin(x)\cos^2(x)-\sin^3x}\) \(\large\color{black}{ \sin(2x+x)=3\sin(x)(1-\sin^2x)-\sin^3x}\) \(\large\color{black}{ \sin(2x+x)=3\sin(x)-3\sin^3x-\sin^3x}\) \(\large\color{black}{ \sin(2x+x)=3\sin(x)-4\sin^3x}\) I don't think they want you to do this, don't they? just approximate it?

OpenStudy (solomonzelman):

it would be too hard to do all of this work....

OpenStudy (perl):

i think we solve that equation sin(3x) = 1/2 sqrt(3)

OpenStudy (perl):

3x = arcsin( sqrt(3) / 2 ) 3x = {pi/3 + 2pi * n , 4pi/3 + 2pi/n }

OpenStudy (solomonzelman):

ohh, I haven't even thought that it is the same as sqrt{3}/2. have I seen sqrt{3}/2... lol